Divide 0001010001001001 (BCD) by 10012 and express the result in octal equivalent.
1
Expert's answer
2020-04-21T19:12:12-0400

Split the code into 4-bit parts:


"0001\space0100\space 0100\space 1001\space (BCD) = 1449_{10}.""1001_2 =1\cdot2^3+0\cdot2^2+0\cdot2^1+1\cdot2^0=9_{10}."


Divide:


"1449_{10}\/9_{10}=161_{10}."

Convert to octal equivalent.

The remainder will be the last smallest number:


"\bigg\{\frac{161}{8}\bigg\}=1."

The integer part is 20, divide by 8:


"\bigg\{\frac{20}{8}\bigg\}=4."

The integer part is 2, divide by 8 (does not have integer part). We have 2:


"\bigg\{\frac{2}{8}\bigg\}=2."


Now join all our numbers from top to bottom, hence:


"161_{10}=241_8."